40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 26 14:54:32 2014, Koji, Summary, LSC, PRMIsb locked with REFL165I&Q again 
    Reply  Wed Mar 26 21:51:42 2014, ericq, Summary, LSC, PRMIsb locked with REFL165I&Q again 
       Reply  Fri Mar 28 17:22:55 2014, Koji, Summary, LSC, PRMIsb locked with REFL165I&Q again 
          Reply  Mon Mar 31 13:15:55 2014, manasa, Summary, LSC, Alignment update 
       Reply  Mon Mar 31 17:47:57 2014, ericq, Summary, LSC, MICH sensing oddities in REFL 3F REFL_33_traj.pdfREFL_165_traj.pdf
          Reply  Mon Mar 31 21:23:30 2014, Gabriele, Summary, LSC, MICH sensing oddities in REFL 3F 
Message ID: 9765     Entry time: Mon Mar 31 13:15:55 2014     In reply to: 9758
Author: manasa 
Type: Summary 
Category: LSC 
Subject: Alignment update 

Quote:

While I'm looking at the PRM ASC servo model, I tried to use the current servo filters for the ASC
as Manasa aligned the POP PDs and QPD yesterday. (BTW, I don't find any elog about it)

 Guilty!!

POP path

The POP PD was showing only ~200 counts which was very low compared to what we recollect from earlier PRMI locks (~400 counts). Also, the POP ASC QPD was also not well-aligned.
While holding PRMI lock on REFL55, I aligned POP path  to its PD (maximize POP DC counts) and QPD (centered in pitch and yaw).

X and Y green

The X green totally lost its pointing because of the misaligned PZTs from last week's power failure. This was recovered.
Y arm green alignment was also recovered.

ELOG V3.1.3-